OpenStudy (anonymous):

A city is building a fence around a rectangular playground. If the perimeter of the playground is 86 feet and the length of one side of the fence is 27 feet, what is the width of the playground? A. 43 feet B. 32 feet C. 16 feet D. 59 feet

OpenStudy (igreen):

The formula for the Perimeter of a rectangle is \(P = 2l + 2w\) Plug in what we know: Perimeter = 86 Length = 27 \(P = 2l + 2w\) \(86 = 2(27) + 2w\) Solve for w: Multiply: \(86 = 54 + 2w\) Subtract 54 to both sides: \(32 = 2w\) Divide 2 to both sides: \(w = 16\)

OpenStudy (igreen):

So the width is 16 feet.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @nerdychic101